OpenStudy (anonymous):

Write an equation in standard form, Ax+By=C, of the line through the two points (0,3) and (2,-3).

OpenStudy (ranga):

Using the two points find the slope m. m = (y2 - y1) / (x2 - x1) The equation of a line in slope intercept form is y = mx + b Substitute m found earlier. The point (0, 3) lies on the line. Put x = 0, y = 3 and solve for b. Once you have y = mx + b, just rearrange it so that it looks like the standard form Ax + By = C